HUBBLE SPACE TELESCOPE DAILY REPORT #2140 PERIOD COVERED: 0000Z (UTC) 06/08/98 - 0000Z (UTC) 06/09/98 Daily Status Report as of 160/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxy NGC-2942 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Five Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. During the final iteration of this proposal, as documented in HSTAR 6519, there were losses of lock that affected one observation (see 2.1 below). Otherwise, the observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1/2/3 7872 (A Morphological Census of z > 1 Cluster Galaxies in the Optical Rest-Frame) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the galactic cluster containing 3C210. Our existing WF/PC-2 imaging of this cluster has let us study the rest-frame ultraviolet light from these galaxies. The proposed NICMOS imaging will allow us to morphologically classify the cluster galaxies at rest-frame optical wavelengths and to place them into the context of the Hubble Sequence. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. During the final iteration of this proposal, as documented in HSTAR 6521, there was one loss of lock affecting one observation. Otherwise, The observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 8051 (STIS Out-of-Focus PSF) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard source GRW+70D582 in order to study the out-of-focus point spread function. As described in HSTAR 6517 and the previous daily report, the guide star re-acquisition for this proposal failed, affecting two observations. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.7 Completed WF/PC-2 8049 (Long WF/PC-2 Decontaminations during 2nd NIC/3 Campaign) The WF/PC-2 was used to make long decontamination procedures as part of the second NIC/3 campaign. Each decon is followed by five 30-minute darks. During this time, no external science WF/PC-2 observations were executed. The decontaminations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/1/2/3 7871 (The Properties of High Redshift Galaxies : A Near-Infrared Redshift Survey) The Near Infrared Camera (Cameras One, Two and Three) was used to make infrared galactic survey observations of high redshifted galaxies in s sky field designated as Field-1417. A comparison between the population of galaxies in the low {z<1} and high {z>2} redshift universe has shown that the peak era of star forming activity in the universe occurs in the redshift interval 1 1 Cluster Galaxies in the Optical Rest-Frame), both of which occur during the period of tomorrow's report, were affected by these losses of lock. 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Operations Notes: ************************************************************* ************************************************************* Over the past two weeks, three HSTARs (6509, 6510, and 6516) have been written on FHST maps using FHST-1. Analysis reveals that all three maps were at the same attitude - one in which the FOV was directed in the vicinity of the galactic plane. This resulted in higher background light during the maps, making it difficult for the FHST to track any suitable stars. This analysis has shown nominal performance of the hardware. In the future, FHST-2 and -3 will be used for all maps and updates at this attitude. During the same period, numerous HSTARs have been written over the past week concerning multiple losses of lock, and failed acquisitions and re-acquisitions (see HSTARs 6519-6522 described in 2.1 above plus HSTARs 6511, 6512, 6514, 6517, and 6518). All of these HSTARs involve a common guide star (star ID 0347800400) using FGS-2. Investigation finds this star to be a binary or double star for FGS-2. This has resulted in failure to achieve fine lock in some cases, and in marginal lock in other cases where sometimes small T3 slews occurred which resulted in immediate losses of lock. FGS 2 performance has been checked, and has been found to be entirely nominal. In addition, the schedule for the next two weeks has been checked and there are two more cases of this particular star/FGS-2 combination (at 160/0702Z and at 160/1024Z). ************************************************************* ************************************************************* Using ROP SR-1, the SSR EDAC error counter was cleared four times. A Slot 1 GENSLEW (162-14A) was performed at 159/2053Z as directed by an operations request. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH